### v1

@@ 3.2 $\mu$LO
Figure 1: Pre-activations behave harmoniously in $\mu$P for LOs and Adam alike. We report the evolution of coordinate-wise standard deviation of the difference between the initial ($t=0$) and $t$-th second-layer pre-activations of an MLP during training (the remaining layers behave similarly, see Sec. G). We observe that all models parameterized in $\mu$P enjoy stable coordinates across scales, while the pre-activations of models in SP blow up after a number of training steps.
@@ 4 Empirical evaluation
We use a suite of optimization tasks to evaluate meta-generalization properties of our $\mu$LOs vs tuned $\mu$Adam [43] and baseline LOs, including pretrained VeLO [24] (the strongest publicly available pre-trained learned optimizer). We focus on evaluating generalization to wider networks, however, we also verify if $\mu$LOs can improve generalization to longer training horizons and deeper networks.
@@
Identifier Type Architecture Optimizee Par. Meta-Training/Tuning Task(s) $\mu\text{LO}_{S}$ Ours small_fc_lopt [23] Adam $\mu$P ImageNet classification, 3-Layer MLP, width $\in\{128\}$ $\mu\text{LO}_{M}$ Ours small_fc_lopt [23] Adam $\mu$P ImageNet classification, 3-Layer MLP, width $\in\{128,512,1024\}$ $\mu$AdamS Baseline Adam $\mu$P ImageNet classification, 3-Layer MLP, width $\in\{128\}$ $\mu$AdamM Baseline Adam $\mu$P ImageNet classification, 3-Layer MLP, width $\in\{1024\}$ $\text{LO}_{S}$ Baseline small_fc_lopt [23] SP ImageNet classification, 3-Layer MLP, width $\in\{128\}$ $\text{LO}_{M}$ Baseline small_fc_lopt [23] SP ImageNet classification, 3-Layer MLP, width $\in\{128,512,1024\}$ VeLO SOTA Baseline VeLO [24] SP We refer the reader to [24, Appendix C.2]
@@
Baseline LOs and $\mu$LOs. The training configuration of each learned optimizer is summarized in Table 1. Each LO is a 2-layer width $32$ per-parameter MLP [23]. The first baseline optimizer, $\text{LO}_{S}$, is trained on a single task where the optimizee is initialized via SP (standard parametrization as in previous works). This is the single task baseline which we assume to meta-generalize poorly to many tasks in our suite. The second baseline optimizer, $\text{LO}_{M}$, is trained on three versions of the same task where the width of the optimizee MLP changes. This baseline sheds light on whether simply varying the optimizee width during meta-training is enough to achieve generalization of the LO to wider networks. Our $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$ are trained on the same tasks as $\text{LO}_{S}$ and $\text{LO}_{M}$, but the optimizee MLPs and parameter updates are parameterized in $\mu$P (section 3.2). For the inner optimization task, we used a maximum unroll length of $1000$ iterations. For all meta-training and hyperparameter tuning details, including their ablation experiments see section D of the appendix.
Pre-trained VeLO. VeLO [24] is a learned optimizer that was meta-trained on a curriculum of progressively more expensive meta-training tasks for a total of $4000$ TPU months. These tasks include but are not limited to image classification with MLPs, ViTs, ConvNets, and ResNets; compression with MLP auto-encoders; generative modeling with VAEs; and language modeling with transformers and recurrent neural networks. During meta-training, VeLO unrolls inner problems for up to 20k steps (20×20$\times$20 × ours); the final model was then fine-tuned on tasks with up to 200k steps of optimization. VeLO, therefore, represents the strongest baseline in our empirical evaluation.
Is Pre-trained VeLO a fair baseline? While we believe the comparison is important given the relevance of our results to scaling up LOs, we highlight that the comparison will unfairly advantage VeLO as all tasks in our suite fall within its meta-training distribution and VeLO was meta-trained on inner unroll horizons well beyond those we evaluate. Thus when comparing our LOs to VeLO, it is important to keep in mind that ours are meta-trained with only $0.004\%$ of VeLO’s compute budget.
@@
$\mu$AdamS and $\mu$AdamM baselines. We tune two Adam baselines that follow $\mu$P Adam [43] on the same meta-training tasks as our learned optimizers and transfer the hyperparameters to the evaluation tasks making them fair baselines (Table 1). We tune the learning rate and three multipliers: input multiplier, output multiplier, and the hidden learning rate multiplier. These multipliers correspond to adding a tunable constant to the pre-activation multiplier for input weights, the pre-activation multiplier for output weights, and the Adam LR for hidden weights. More details about the grid search over $500$ configurations are provided in Section C of the appendix.
@@
Evaluation tasks. Our evaluation suite includes 20 tasks spanning image classification (CIFAR-10, ImageNet) using MLPs and Vision Transformers (ViTs) [12] and autoregressive language modeling with a decoder-only transformer on LM1B [8]. To create the tasks, we further vary image size (for image classification), width and depth of the optimizee network, and the number of optimization steps. See Table 4 of the appendix for an extended description of all the tasks.
((a))
((b))
((c))
Figure 2: Evaluating generalization to wider networks. We train MLPs of different sizes to classify $32\times 32\times 3$ ImageNet images. Our optimizers are meta-trained for $1000$ inner steps (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that $\mu\text{LO}_{M}$ generalize smoothly to all tasks, even surpassing the performance of VeLO (Fig.(b),(c)). We plot average training loss over $5$ seeds with standard error bars.
((a))
((b))
((c))
Figure 3: Meta-generalization to larger images and wider networks. We train MLPs of different sizes to classify $64\times 64\times 3$ ImageNet images. Our optimizers are meta-trained for $1000$ inner steps (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that $\mu\text{LO}_{M}$ generalizes smoothly to all tasks, obtaining the best performance on the largest task. We plot average training loss over $5$ seeds with standard error bars.
4.1 Results
In the following sections, we first (section 4.1.1) present results empirically verifying the pre-activation stability of our $\mu$LOs and subsequently present the results of our empirical evaluation of meta-generalization to wider networks, deeper networks, and longer training horizons. In section 4.1.2, we start by evaluating our learned optimizers and the corresponding baselines in-distribution and on wider versions of the meta-training task(s) (Figure 2). Continuing with our evaluation of large widths, we study generalization to larger input images (e.g., a wider input layer) and wider hidden layers in Figure 3 and generalization to Cifar-10 (e.g., a smaller output layer) and wider hidden layers in Figure 4. To conclude our study of generalization to larger width, in section 4.1.3, we evaluate our optimizers on out-of-distribution ViT image classification and autoregressive language modeling (figure 5). In section 4.1.4, we transition to evaluating deeper versions of the MLP, ViT, and LM tasks seen previously (see figure 6). Finally, in section 4.1.5, we evaluate generalization to longer unrolls (25$\times$ meta-training) for MLP, ViT, and LM tasks (see figure 7).
4.1.1 Evaluating pre-activation stability
We now verify that desiderata J.1 of [43] is satisfied empirically. In Figure 1, we report the evolution of the coordinate-wise standard deviation of the difference between initial (t=0) and current (t) second-layer pre-activations of an MLP during the first $500$ steps of training. We observe that all models parameterized in $\mu$P enjoy stable coordinates across widths, while the pre-activations of the larger models in SP blow up after a number of training steps. Notably, SP Adam’s pre-activations blow up immediately, while $\text{LO}_{S}$ and $\text{LO}_{M}$ take longer to blow up and have a more erratic pattern; we hypothesize that this is a side effect of meta-training where the optimizers may learn to keep pre-activations small by rescaling updates. Section G of the appendix contains similar plots for the remaining layers of the MLP which show similar trends.
((a))
((b))
((c))
Figure 4: Meta-generalization to Cifar-10 and wider networks. We train MLPs of different sizes to classify Cifar-10 images. Our optimizers are meta-trained for $1000$ inner steps (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that $\mu\text{LO}_{M}$ generalizes seamlessly to all tasks, reaching 00 loss before VeLO in subplots (b) and (c) while all other optimizers struggle to train stably. We plot average training loss over $5$ seeds with standard error bars.
4.1.2 Meta-generalization to wider networks
Wider hidden layers. Figure 2 reports the results for optimizers evaluated on tasks ${\mathcal{T}}^{\text{MLP}}_{1},{\mathcal{T}}^{\text{MLP}}_{3},{\mathcal{T}}^{\text{MLP}}_{4}$, which involve training an MLP of width W=128, W=2048 and W=8192 (at least 8 times wider compared to those used in meta-training). The MLPs are trained on $32\times 32\times 3$ images of ImageNet for $5000$ steps. Since LOs are meta-trained for $1000$ inner steps, steps $1000-5000$ are considered out-of-distribution, posing an additional challenge for meta-generalization. Among the optimizers, $\text{LO}_{S}$ performs the worst diverging after around 1000 steps (Fig. 2, (a)) or immediately (Fig. 2, (b) and (c)). In contrast, our $\mu\text{LO}_{S}$ generalizes seamlessly to the larger tasks matching (Fig. 2, (b)) or improving upon VeLO (Fig. 2, (c)). However, $\mu\text{LO}_{S}$ diverges on the small task (Fig. 2, (a)), which may be caused by slight overfitting on the meta-training task leading to larger updates causing divergence. Compared to $\mu\text{LO}_{S}$, our $\mu\text{LO}_{M}$ generalizes seamlessly without diverging. This generalization is especially notable as its counterpart in SP, $\text{LO}_{M}$, comes nowhere close. Moreover, $\mu\text{LO}_{M}$ beats VeLO (Fig. 2, (b) and (c)) despite VeLO’s massive meta-training budget. While $\mu$Adam also shows strong performance, it underperforms our $\mu\text{LO}_{M}$ in all cases.
Larger images and wider hidden layers. Figure 3 reports the results for LOs evaluated on tasks ${\mathcal{T}}^{\text{MLP}}_{5}$-${\mathcal{T}}^{\text{MLP}}_{7}$, which involve training an MLP of width W=128, W=512 and W=2048 on $64\times 64\times 3$ images of ImageNet. This task evaluates generalization to a wider input layer and, therefore, all tasks are considered out-of-distribution. The $\text{LO}_{S}$ baseline is incapable of generalizing to these tasks, diverging at the beginning in each case. $\text{LO}_{M}$ performs slightly better indicating that meta-training on more tasks helps to some extent, however it still converges poorly for W=128 and W=512 and diverges for W=2048. Unlike $\text{LO}_{S}$, our $\mu\text{LO}_{S}$ only diverges after $1000$ steps of optimization and improves its performance for longer unrolls as the task size increases (similarly to Fig. 2). Once again, $\mu\text{LO}_{M}$ generalizes seamlessly to all tasks without encountering any instabilities outperforming VeLO and other optimizers at W=2048. The $\mu$Adam baselines generalize well to these tasks, however as in previous experiments, they underperform our $\mu\text{LO}_{M}$ overall.
CIFAR-10 and wider hidden layers. Figure 4 reports the results of optimizers evaluated on tasks ${\mathcal{T}}^{\text{MLP}}_{8}$,${\mathcal{T}}^{\text{MLP}}_{10}$, and ${\mathcal{T}}^{\text{MLP}}_{11}$ (see ${\mathcal{T}}^{\text{MLP}}_{9}$ results in Fig. 10 of the appendix). Each of these tasks involves training an MLP of different width to classify CIFAR-10 images. These tasks can be seen as evaluating generalization to a smaller output layer size while width is increased. Aligned with our previous experiments, the baseline $\text{LO}_{S}$ and $\text{LO}_{M}$ struggle to generalize either diverging or converging slowly. Our $\mu\text{LO}_{S}$ is more stable, however $\mu\text{LO}_{M}$ is again the best on larger networks monotonically decreasing the loss. Notably, it even reaches loss=0 before VeLO at width W=512,2048,8192. In this case, we observe much superior generalization on the part of $\mu\text{LO}_{M}$ when compared to $\mu$AdamS and $\mu$AdamM whose training curves are fraught with instability.
In summary, the results in Fig. 2,3 and 4 indicate that simply meta-training on more tasks is not enough to achieve strong meta-generalization, while $\mu$LO can enable such meta-generalization even outperforming VeLO in some cases.
@@ 4.1.2 Meta-generalization to wider networks
((f))
Figure 5: Meta-generalization to wider out-of-distribution tasks. We report the performance of our optimizers on autoregressive language modeling (top) and ViT image classification (bottom). The datasets used are LM1B and $32\times 32\times 3$ ImageNet, respectively. We observe that $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$ train stably and consistently outperform all optimizers except VeLO (LM) our outperform SP LOs and match $\mu$Adam (ViT). We plot average training loss over $5$ seeds with standard error bars.
4.1.3 Meta-generalization far out-of-distribution
In the previous section, some of the tasks were out-of-distribution but remained similar. In this sections, we evaluate on the drastically different tasks than those used to meta-train our optimizers.
Language modeling. We train 3-layer decoder-only transformers of different widths on LM1B for $5000$ steps. The batch size is set to $128$ and the sequence length is set to $64$. Figure 5 (a,b,c) reports the training loss over iterations for each optimizer. We observe smooth training curves for $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$ which best all optimizers except VeLO at every width. In contrast, $\text{LO}_{S}$ and $\text{LO}_{M}$ encounter training instabilities at $w=1024,2048$ and $w=2048$ respectively, significantly underperforming $\mu$LOs. While the $\mu$Adam baselines train stably, they underperform $\mu$LOs. Finally, VeLO performs best of all on all tasks which is expected given its compute budget (more than 28,000×28,000$\times$28 , 000 × the compute used by $\mu\text{LO}_{M}$) and the fact that these tasks (including batch size) fall under its meta-training distribution.
ViT image classification. We train 3-layer ViTs of different widths to classify $32\times 32\times 3$ ImageNet images. The batch size is set to $1024$ Figure 5 (d,e,f) reports the training loss over iterations for each optimizer. We observe smooth training curves for $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$ which best SP LOs at each width and improve on $\mu$Adam as width increases. While $\text{LO}_{S}$ and $\text{LO}_{M}$ do not diverge, for larger models they have difficulty decreasing loss after the first $500$ steps and they significantly underperform $\mu$ LOs across the board. Similarly to the language task, VeLO performs the best as expected.
In summary, even for out-of-distribution tasks, we observe seamless generalization for our $\mu$LOs, showing that using $\mu$P can significantly improve optimizer quality without increasing meta-training time.
4.1.4 Meta-generalization to deeper networks
In this section, we evaluate LO meta-generalization to deeper networks. Specifically, we increase the number of layers used in MLP, ViT, and LM tasks from $3$ to $16$, while being sure to select models that have widths within the meta-training range ($128-1024$) to avoid confounding the results. Figure 6 reports the performance of our multi-task learned optimizers on deeper networks. We observe that $\mu\text{LO}_{M}$ optimizes stably throughout and outperforms $\text{LO}_{M}$ by the end of training on each task, despite being meta-trained on MLPs of exactly the same depth. Moreover, $\text{LO}_{M}$ immediately diverges when optimizing the deep MLP while $\mu\text{LO}_{M}$ experience no instability. This is remarkable as, unlike width, there is no theoretical justification for $\mu$P’s benefit to deeper networks. We hypothesizes that $\mu$P’s stabilizing effect on networks activations leads to this improvement generalization.
In summary, we find, empirically, that using $\mu$P during meta-training benefits the generalization of learned optimizers to deeper networks.

### v2

@@ 3.3 $\mu$LO Meta-training Recipe
Figure 3: Pre-activations behave harmoniously in $\mu$P for LOs and Adam alike. We report the evolution of coordinate-wise standard deviation of the difference between the initial ($t=0$) and $t$-th second-layer pre-activations of an MLP during training (the remaining layers behave similarly, see Sec. F). We observe that all models parameterized in $\mu$P enjoy stable coordinates across scales, while the pre-activations of models in SP blow up after a number of training steps.
@@ 4 Empirical evaluation
We use a suite of optimization tasks of varying width to evaluate meta-generalization properties of our $\mu$LOs vs tuned $\mu$Adam (Yang et al., 2022), SP AdamW Loshchilov and Hutter (2019), and baseline SP LOs. We also include pre-trained VeLO (Metz et al., 2022b) as an oracle which we denote as VeLO-4000. Meta-trained for $4000$ TPUv4 months, it is the strongest publicly available pre-trained learned optimizer. We focus on evaluating generalization to wider networks, however, we also verify whether $\mu$LOs can improve generalization to longer training horizons and deeper networks.
@@
Identifier Type Architecture Optimizee Par. Meta-Training / Tuning Task(s) $\mu\text{LO}_{S}$ Ours small_fc_lopt (Metz et al., 2022a) $\mu$LO Sec. 3.2 ImageNet classification, 3-Layer MLP, width $\in\{128\}$ $\mu\text{LO}_{M}$ Ours small_fc_lopt (Metz et al., 2022a) $\mu$LO Sec. 3.2 ImageNet classification, 3-Layer MLP, width $\in\{128,512,1024\}$ $\mu$VeLOM Ours VeLO (Metz et al., 2022b) $\mu$LO Sec. 3.2 ImageNet classification, 3-Layer MLP, width $\in\{128,512,1024\}$ $\text{LO}_{S}$ LO Baseline small_fc_lopt (Metz et al., 2022a) SP ImageNet classification, 3-Layer MLP, width $\in\{128\}$ $\text{LO}_{M}$ LO Baseline small_fc_lopt (Metz et al., 2022a) SP ImageNet classification, 3-Layer MLP, width $\in\{128,512,1024\}$ VeLOM LO Baseline VeLO (Metz et al., 2022b) SP ImageNet classification, 3-Layer MLP, width $\in\{128,512,1024\}$ VeLO-4000 Oracle LO Baseline VeLO (Metz et al., 2022b) SP We refer the reader to (Metz et al., 2022b, Appendix C.2) $\mu$AdamM Baseline $\mu$P Adam ImageNet classification, 3-Layer MLP, width $\in\{1024\}$ AdamW Baseline SP ImageNet classification, 3-Layer MLP, width $\in\{1024\}$
@@
Baseline LOs and $\mu$LOs. The meta-training configuration of each learned optimizer is summarized in Table 1. Each learned optimizer (ours and the baselines) in our empirical evaluation is meta-trained using the multiple-width single-task meta-training recipe proposed in section 3.3. The baseline sheds light on whether simply varying the SP optimizee width during meta-training is enough to achieve generalization of the LO to wider networks in SP. During meta-training, we set the inner problem length to be $1000$ iterations. Therefore, any optimization beyond this length is considered out-of-distribution. For all meta-training and hyperparameter tuning details, including ablation experiments, see section C of the appendix.
@@
$\mu$Adam $\mu$Adam is a strong hand-designed $\mu$P baseline. It follows the Yang et al. (2022) Adam $\mu$-parametrization and is tuned on the largest meta-training task seen by our learned optimizers (Table 1). We tune the learning rate and three multipliers: input multiplier, output multiplier, and the hidden learning rate multiplier. These multipliers correspond to adding a tunable constant to the pre-activation multiplier for input weights, the pre-activation multiplier for output weights, and the Adam LR for hidden weights. More details about the grid search over $500$ configurations are provided in Section B.1 of the appendix.
AdamW AdamW (Loshchilov and Hutter, 2019) is a strong hand-designed SP baseline. It is tuned on the largest meta-training task seen by our learned optimizers (Table 1). We tune the learning rate, $\beta_{1}$,$\beta_{1}$, and the weight decay. More details about the grid search over $500$ configurations are provided in Section B.1 of the appendix.
Pre-trained VeLO (VeLO-4000). VeLO (Metz et al., 2022b) is a learned optimizer that was meta-trained on a curriculum of progressively more expensive meta-training tasks for a total of $4000$ TPU months. These tasks include but are not limited to image classification with MLPs, ViTs, ConvNets, and ResNets; compression with MLP auto-encoders; generative modeling with VAEs; and language modeling with transformers and recurrent neural networks. During meta-training, VeLO-4000 unrolls inner problems for up to 20k steps (20×20$\times$20 × ours); the final model was then fine-tuned on tasks with up to 200k steps of optimization. VeLO-4000, therefore, represents the strongest baseline in our empirical evaluation and we consider it to be an oracle.
Is VeLO-4000 a fair baseline? While we believe the comparison is important given the relevance of our results to scaling up LOs, we highlight that the comparison will unfairly advantage VeLO-4000 as all tasks in our suite fall within its meta-training distribution and VeLO-4000 was meta-trained on inner unroll horizons well beyond those we evaluate. Thus, when comparing our LOs to VeLO-4000, it is important to keep in mind that ours are meta-trained with only $0.004\%$ of VeLO-4000’s compute budget.
@@
Evaluation tasks. Our evaluation suite includes $35$ tasks spanning image classification (CIFAR-10, ImageNet) using MLPs and Vision Transformers (ViTs) (Dosovitskiy et al., 2020) and autoregressive language modeling with a decoder-only transformer on LM1B (Chelba et al., 2013). To create the tasks, we further vary image size (for image classification), width, and depth of the optimizee network, and the number of optimization steps. See Table 7 of the appendix for an extended description of all the tasks.
@@
Figure 4: Evaluating generalization to wider networks for different tasks. Tasks Our optimizers are meta-trained for $1000$ inner steps (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We plot average training loss over $5$ seeds with standard error bars. We observe that $\mu\text{LO}_{M}$ and $\mu$VeLOMgeneralize smoothly to longer unrolls and all unseen tasks, unlike their SP counterparts which diverge or failt to make progress. $\mu$LOs even surpass or match the performance of VeLO in subfigures (a), (b), and (c)). Moreover, they also substantially best the well-tuned hand-designed baselines on LM and ViT tasks (subfigures (d) and (e)) and best or match the best performing hand-designed optimizer in subfigures (a),(b), and (c).
4.1 Results
In the following sections, we first (Sec. 4.1.1) present results empirically verifying the pre-activation stability of our $\mu$LOs. Subsequently, we present the results of our main empirical evaluation of meta-generalization to wider networks (Sec. 4.1.1), a study of $\mu$LOs generalization to deeper networks (Sec. 4.1.3), and a study of $\mu$LOs generalization to longer training horizons (Sec. 4.1.4).
4.1.1 Evaluating pre-activation stability
We now verify that desiderata J.1 of Yang et al. (2022) is satisfied empirically. In Figure 3, we report the evolution of the coordinate-wise standard deviation of the difference between initial (t=0) and current (t) second-layer pre-activations of an MLP during the first $500$ steps of training. We observe that all models parameterized in $\mu$P enjoy stable coordinates across widths, while the pre-activations of the larger models in SP blow up after a number of training steps. Notably, SP Adam’s pre-activations blow up immediately, while $\text{LO}_{S}$ and $\text{LO}_{M}$ take longer to blow up and have a more erratic pattern; we hypothesize that this is a side effect of meta-training where the optimizers may learn to keep pre-activations small by rescaling updates. Section F of the appendix contains similar plots for the remaining layers of the MLP which show similar trends.
Table 2: In-distribution and out-of-distribution average performance of optimizers. We report the average rank of different optimizers across the five tasks in our suite. We evaluate in-distribution at a Base width of $1024$ as this is the width used to tune the hand-designed baselines. We also evaluate out-of-distribution widths: Large (2048) and XL (largest size for each task see Tab.7 of the appendix). We bold the strongest, underline the second strongest, and italicize the third strongest average rank in each column. We do not bold entries of VeLO-4000 as it is reported only for reference since it is not a fair comparison. We observe that, across all iterations, when compared to fair baselines, $\mu\text{LO}_{M}$ obtains the best rank for all settings except for the XL task at $5000$ iterations, where it is only bested by $\mu$VeLO.
Loss at 1k steps Loss at 3k steps Loss at 5k steps Optimizer ID (Base) OoD (Large) OoD (XL) ID (Base) OoD (Large) OoD (XL) ID (Base) OoD (Large) OoD (XL) AdamW 3.40 3.20 4.60 3.60 3.80 4.80 4.00 4.40 4.80 $\mu$Adam 4.40 4.20 4.00 3.60 3.60 3.40 3.60 3.60 3.20 VeLOM 5.00 5.80 5.60 6.80 6.40 7.00 7.00 7.00 6.80 LOM 5.20 6.60 6.80 5.20 6.60 6.00 5.60 6.00 6.20 $\mu$VeLOM (ours) 4.40 3.40 2.20 3.60 2.60 2.20 4.00 2.80 2.00 $\mu$LOM (ours) 2.80 1.80 2.00 2.40 2.00 1.80 2.20 2.00 2.80 VeLO-4000 2.80 3.00 2.80 2.80 3.00 2.80 1.60 2.20 2.20
4.1.2 Meta-generalization to wider networks
Given our goal of improving LO generalization to unseen wider tasks, the bulk of our empirical evaluation is presented in this section. Specifically, we evaluate the behavior of $\mu$LOs as the width of tasks increases well beyond what was seen during meta-training. To accomplish this, we fix the depth of each task and vary the width (see Table 7 for a full list of tasks), leading to a testbed of $32$ different tasks. We then train each task using the baselines and $\mu$-optimizers outlined in section 4 for $5000$ steps for $5$ different random seeds. This involves training $1120$ different neural networks. To make the results easily digestible, we summarize them by width and final performance in Figure 4(c) and by average optimizer rank in Table 2. We also highlight the smooth training dynamics of our optimizers at the largest widths in figure  4(c).
Performance measured by final loss as a function of width
Figure 1 compares the training loss after $1000$ steps of SP learned optimizers to $\mu$-parameterized learned optimizers for different widths. This is shown in three subfigures for three MLP image classification tasks: (a) Imagenet $32\times 32\times 3$ (IN32), (b) Imagenet $64\times 64\times 3$ (IN64), and (c) Cifar-10 $32\times 32\times 3$ (C10). Subfigure (a) shows the performance of learned optimizers on larger versions of the meta-training tasks. We observe that the $\mu$LOs achieve lower final training loss as the width of the task is increased. In contrast, $\text{LO}_{M}$ diverges for widths larger than $2048$ and VeLOM fails to substantially decrease the loss at larger widths, falling behind the $\mu$LOs. Subfigure (b) evaluates our $\mu$LOs of larger ImageNet images (e.g., when the input width is larger). Similarly, we observe smooth improvements in the loss as the optimizee width increases for $\mu$LOs, while their SP counterparts either diverge at width $512$ (LOM) or fail to substantially improve the loss beyond width $1024$ (VeLOM). Finally, Subfigure (c) shows the performance of our $\mu$LOs on Cifar-10 (smaller output width) as the width of the model is increased. Similarly, we observe smooth improvements in the loss as the width increases for $\mu$LOs, while their SP counterparts either diverge immediately at small widths (VeLOM) or diverge by width $1024$ (LOM).
Performance measured by average optimizer rank
Table 2 reports the average rank of different optimizers on in-distribution width tasks (Base, width 1024) and out-of-distribution width tasks (Large (width 2048) and XL (maximum width)). Each entry of the table corresponds to the optimizer’s average rank (within the $7$ optimizers evaluated) over the $5$ tasks in our suite: Cifar 10 MLP image classification, ImageNet 32 MLP image classification, ImageNet 64 MLP image classification, ImageNet 32 ViT image classification, and LM1B transformer language modelling. The optimizers are ranked by their training loss at the given iteration. We report average ranks for $1000$ iterations (inner-problem length), $3000$ iterations, and $5000$ iterations. We bold the strongest, underline the second strongest, and italicize the third strongest average rank in each column. We do not bold entries of VeLO-4000 as it is reported only for reference since it is not a fair comparison. We observe that, across all iterations, when compared to fair baselines, $\mu\text{LO}_{M}$ obtains the best rank for all settings except for the XL task at $5000$ iterations, where it is only bested by $\mu$VeLO. When only looking at the out-of-distribution Large and XL tasks, we observe that $\mu\text{LO}_{M}$ and $\mu$VeLOMdominate the first two spots of the optimizer podium in all cases except one. For the Large task at $1000$ steps, $\mu$VeLOMis bested by AdamW. When comparing our $\mu$LOs to VeLO-4000, we observe that at least one of $\mu\text{LO}_{M}$ and $\mu$VeLOMbests VeLO-4000 on all tasks except for the large task at $5000$ iterations. This is remarkable as our $\mu$LOs are trained on many orders of magnitude less compute than VeLO-4000. These results demonstrate that meta-training LOs using our recipe yields substantial improvements in meta-generalization (across various tasks and widths) over LOs from previous work and strong hand-designed baselines.
Training dynamics at the largest widths
Figure 4(c) reports the training curves of different optimizers on the largest width tasks in our suite. Despite training for 5×5$\times$5 × longer than the maximum meta-training unroll length, our $\mu$LOs are capable of smoothly decreasing the loss for the largest out-of-distribution tasks in our suite. In contrast, the strong SP LO baselines diverge by $1000$ steps (subfigures (a),(b),(c),(d)), or fail to decrease the training loss (subfigure (e)). Our $\mu$LOs also substantially best the well-tuned hand-designed baselines on LM and ViT tasks (subfigures (d) and (e)) and best or match the best performing hand-designed optimizer in subfigures (a),(b), and (c). When comparing with VeLO-4000, we observe that our $\mu$LOs substantially outperform VeLO in subfigures (a),(b), and $\mu\text{LO}_{M}$ outperforms VeLO-4000 in subfigure (c). In contrast, VeLO-4000 outperforms our $\mu$LOs on transformer language modeling and ViT image classification, the most out-of-distribution tasks for them. These findings show that $\mu$LOs can outperform VeLO-4000 on larger in-distribution tasks, suggesting that scaling meta-training in SP (e.g., as done for VeLO) may not be sufficient to achieve strong meta-generaliztion to the largest tasks, but that meta-training in $\mu$P could be.
In summary, the results in Fig. 1,Tab. 2 and Fig.4(c) demonstrate that our $\mu$LO meta-training recipe represents a considerable advancement to low-cost meta-generalization for learned optimizers. The technique is shown to be a substantial improvement over previous work.